Grade 12Algebra

Multiply and reduce to lowest form and convert into a mixed fraction:
13 × (1/3)

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

By the rule Multiplication of fraction,
Product of fraction = (product of numerator)/ (product of denominator)
Then,
= (13/1) × (1/3)
= (13 × 1)/ (1 × 3)
= (13/3)
= 4 whole 1/3

Thanks